OpenStudy (oswaldmurphy):

Choose the equation below that represents the line passing through the point (2, – 5) with a slope of −3. I also need help using a graphing calculator and which is the best online one to use?

pooja195 (pooja195):

Hi Welcome to OpenStudy :) Online graphing calculator to use: https://www.desmos.com/calculator And please post your options! :)

OpenStudy (oswaldmurphy):

y = −3x − 13 y = −3x + 11 y = −3x + 13 y = −3x + 1 I have been using desmos but I don't understand how. Thanks!

pooja195 (pooja195):

y=mx+b m=slope b=y-intercept What have you done so far?

OpenStudy (oswaldmurphy):

\[y = -3x + \frac{ 2 }{ -5 }\]

TheSmartOne (thesmartone):

Pooja had to leave, so she asked me to help you finish up. :)

OpenStudy (oswaldmurphy):

What do I do next? I posted the equation. Is that correct from the information I was given?

TheSmartOne (thesmartone):

ok, I see you got the slope correct, but you didn't get the y-intercept correct. You were given the point (2, -5) to find the y-intercept, you do not make it 2/-5 To find the y-intercept you plug it back in to the equation so we have y = mx + b we know m which is the slope is -3. What do we have now? y = -3x + b b is the y-intercept and is what we're trying to find :) now plug in the point. (2, -5) is in the form of (x, y) so that means plug in x = 2 and y = -5 and solve for b, also known as the y-intercept :)

OpenStudy (oswaldmurphy):

So -5 = -6 + b?

TheSmartOne (thesmartone):

correct, now add 6 on both sides to solve for b :)

OpenStudy (oswaldmurphy):

Ok, how do I use a graphing calculator?

TheSmartOne (thesmartone):

to use a graphing calculator for this question, I would plug in the point they gave you and then plug in the answer choices to see which equation goes through the point

OpenStudy (oswaldmurphy):

How? I have never even seen one in my life :P

TheSmartOne (thesmartone):

https://www.desmos.com/calculator/ktvrpmgcyi

OpenStudy (oswaldmurphy):

Thank you!

TheSmartOne (thesmartone):

Anytime! :)